OpenStudy (anonymous):

The screen of a rectangular television set is 5 inches wider than it is high. If the perimeter of the screen is 74 inches, find its dimensions.

OpenStudy (kropot72):

Let height = h Let w = width 74 = 2h + 2w w = h + 5 Substitute the value for w in the first equation 74 = 2h + 2(h + 5) 74 = 2h + 2h + 10 4h = 74 - 10 = 64 Can you find the values of h and w now?

OpenStudy (kropot72):

4h = 64 h = 64/4 = 16 inches w = 16 + 5 = 21 inches